# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 2.3   Present status
The efforts at the various laboratories discussed above have significantly increased the number of known excitations of the nucleon and their decay modes. A survey of the known light-quark baryons including hyperons is shown in Fig. 2.2. Table 2.3 compares the current status with the status as of 20 years ago [485]. In the mass region from 1850 MeV to 2150 MeV, we now recognize 11 $N^{*}$ resonances, compared to four known in 2004. These 11 resonances have a total overall star rating of 30 stars; in 2004, they had 7 overall stars. The star rating was attributed to their $N\pi$ coupling, with only 11 additional stars given for non-$N\pi$ decay modes, all with one-star evidence. Currently, for resonances in the mass range from 1850 MeV to 2150 MeV, 107 stars are attributed to non-$N\pi$ decay modes. Our knowledge of high-mass resonances comes primarily from elastic scattering of $\pi N$. Table 2.2 lists their PDG masses, widths and uncertainties given in the RPP. For 1* and 2* resonances, the uncertainties are from the measurements; for multiple measurements, the uncertainties span the range of reported data.
Table 2.3: $N^{\ast}$ and $\Delta^{*}$ resonances. Pole positions in terms of the mass $M$ and width $\Gamma$ (in MeV) are taken from Ref. [474, 475]. Their decay channels as of 2004 [485] are shown in black and those from 2024 in red and in brown. Recent findings are highlighted in blue. Blue stars are suggestions by us. A candidate reported as observed in electroproduction is shown in green. Resonances in cyan are not seen in Ref. [474, 475]. Their pole positions are from Ref. [3]. We use the symbol $\sigma$ for the $f_{0}(500)$ throughout this paper.
@@
Figure 2.3: (Color online) Modeling the transition of the universe from the quark-gluon phase to the hadron phase during the first microseconds after the Big Bang. Left: Ratio of strangeness to baryon chemical potential near the critical point in the transition. The shaded area shows the projections from Lattice QCD. The dashed lines represent the hadron resonance gas model projections with all known hadrons as cataloged by the PDG in the RPP editions in 2012 (red dashed-dotted) and 2016 (green dashed-dotted), including ∗∗∗ and ∗∗∗∗ states. Right: The $N^{*}$ and $\Delta^{*}$ resonances are the additional 2016 states with only and ∗∗ states, which are unconfirmed. If a number of those could be confirmed experimentally they could have a significant impact on a more complete understanding of this fundamental hadron phase transition, as the black dashed line indicates. The graphics is adapted from  [486].
@@ 2.4   Electroproduction of pseudoscalar mesons
The analysis tools for pion and photoproduction of mesons off nucleons are described in detail in recent review articles, in particular in [14]. The formalism used to analyze the electroproduction data is accessible only in the original articles. Therefore, the methods used are described in some detail in this review.
@@ 2.4.1   Formalism
The kinematics for single $\pi^{+}$ production is shown in Fig. 2.4. The observables of the process $\gamma_{v}p\to\pi N^{\prime}$ can be expressed in terms of six parity-conserving helicity amplitudes [15, 487, 488]
${H_{i}=\left<\lambda_{\pi};\lambda_{N}|T|\lambda_{\gamma_{v}};\lambda_{p}\right>,}\vspace{-2mm}$ (2.6)
where $\lambda$ denotes the helicity of the respective particle, $\lambda_{\pi}=0$, $\lambda_{N}=\pm\frac{1}{2}$, $\lambda_{\gamma_{v}}=\pm 1,0$, and $\lambda_{p}=\pm\frac{1}{2}$, and $H_{i}$ are complex functions of $Q^{2}$, $W$, and $\theta^{*}_{\pi}$.
@@
$\begin{array}[]{rl}\sigma_{T}&\!\!\!=\displaystyle\frac{|{\mbox{\boldmath$p$}}_{\pi}|W}{2KM}\,(|H_{1}|^{2}+|H_{2}|^{2}+|H_{3}|^{2}+|H_{4}|^{2})\,,\\[11.38109pt] \sigma_{TT}&\!\!\!=\displaystyle\frac{|{\mbox{\boldmath$p$}}_{\pi}|W}{2KM}\,\text{Re}\,(H_{2}H_{3}^{*}-H_{1}H_{4}^{*})\,,\end{array}\qquad\begin{array}[]{rl}\sigma_{L}&\!\!\!=\displaystyle\frac{|{\mbox{\boldmath$p$}}_{\pi}|W}{2KM}\,(|H_{5}|^{2}+|H_{6}|^{2})\,,\\[11.38109pt] \sigma_{LT}&\!\!\!=\displaystyle\frac{|{\mbox{\boldmath$p$}}_{\pi}|W}{2KM}\,\text{Re}\,[H_{5}^{*}(H_{1}-H_{4})+H_{6}^{*}(H_{2}+H_{3})]\,,\end{array}$ (2.7)
@@
$\begin{array}[]{rl}H_{1}&\!\!\!=\displaystyle\frac{1}{\sqrt{2}}\sin\theta\cos{\frac{\theta}{2}}\sum_{l=1}^{\infty}(B_{l+}-B_{(l+1)-})(P^{\prime\prime}_{l}-P^{\prime\prime}_{l+1})\,,\\[11.38109pt] H_{2}&\!\!\!=\displaystyle\sqrt{2}\cos{\frac{\theta}{2}}\sum_{l=1}^{\infty}(A_{l+}-A_{(l+1)-})(P^{\prime}_{l}-P^{\prime}_{l+1})\,,\\[11.38109pt] H_{3}&\!\!\!=\displaystyle\frac{1}{\sqrt{2}}\sin\theta\sin{\frac{\theta}{2}}\sum_{l=1}^{\infty}(B_{l+}+B_{(l+1)-})(P^{\prime\prime}_{l}+P^{\prime\prime}_{l+1})\,,\end{array}\qquad\begin{array}[]{rl}H_{4}&\!\!\!=\displaystyle\sqrt{2}\sin{\frac{\theta}{2}}\sum_{l=1}^{\infty}(A_{l+}+A_{(l+1)-})(P^{\prime}_{l}+P^{\prime}_{l+1})\,,\\[11.38109pt] H_{5}&\!\!\!=\displaystyle\sqrt{2}\cos{\frac{\theta}{2}}\sum_{l=1}^{\infty}(C_{l+}-C_{(l+1)-})(P^{\prime}_{l}-P^{\prime}_{l+1})\,,\\[11.38109pt] H_{6}&\!\!\!=\displaystyle\sqrt{2}\sin{\frac{\theta}{2}}\sum_{l=1}^{\infty}(C_{l+}+C_{(l+1)-})(P^{\prime}_{l}+P^{\prime}_{l+1})\,,\end{array}$ (2.9)
@@
$\begin{array}[]{rl}M_{l+}&\!\!\!=\displaystyle\frac{1}{2(l+1)}\,[2A_{l+}-(l+2)B_{l+}]\,,\\[11.38109pt] M_{l+1,-}&\!\!\!=\displaystyle\frac{1}{2(l+1)}\,(2A_{l+1,-}-lB_{l+1,-1})\,,\\[11.38109pt] S_{l+}&\!\!\!=\displaystyle\frac{1}{l+1}\sqrt{\frac{{{\mbox{\boldmath$Q^{*}$}}^{2}}}{Q^{2}}}\,C_{l+}\,,\end{array}\qquad\begin{array}[]{rl}E_{l+}&\!\!\!=\displaystyle\frac{1}{2(l+1)}\,(2A_{l+}+lB_{l+})\,,\\[11.38109pt] E_{l+1,-}&\!\!\!=\displaystyle\frac{1}{2(l+1)}\,[-2A_{l+1,-}+(l+2)B_{l+1,-}]\,,\\[11.38109pt] S_{l+1,-}&\!\!\!=\displaystyle\frac{1}{l+1}\sqrt{\frac{{{\mbox{\boldmath$Q^{*}$}}^{2}}}{Q^{2}}}\,C_{l+1,-}\,,\end{array}$ (2.10)
@@
$\begin{array}[]{rl}\hat{A}_{l\pm}&\!\!\!=\displaystyle\mp F\,C^{I}_{\pi N}\,A_{1/2}\,,\\[11.38109pt] \hat{S}_{l\pm}&\!\!\!=\displaystyle-F\,\frac{2\sqrt{2}}{2J+1}\,C^{I}_{\pi N}\,S_{1/2}\,,\end{array}\qquad\begin{array}[]{rl}\hat{B}_{l\pm}&\!\!\!=\displaystyle\pm F\,\sqrt{\frac{16}{(2j-1)(2j+3)}}\,C^{I}_{\pi N}\,A_{3/2}\,,\\[11.38109pt] F&\!\!\!=\displaystyle\sqrt{\frac{1}{(2j+1)}\pi}\,\frac{K}{p_{\pi}}\,\frac{\Gamma_{\pi}}{\Gamma^{2}}\,,\end{array}$ (2.11)
where the $C^{I}_{\pi N}$ are isospin coefficients. The subscripts $1/2$ and $3/2$ represent the total helicity ±$\pm 1/2$ and ±$\pm 3/2$ of the photon-proton system $\gamma^{*}p$ in the initial state. The total transverse absorption cross section for the transition into a specific resonance is given by
@@
and similarly for the longitudinal (scalar) total cross section $\sigma_{L}$. The quantities $A_{1/2},~{}A_{3/2},~{}S_{1/2}$ are the transverse and scalar transition amplitudes for resonances excited through photon-proton interactions. Resonances with $J={1/2}$ have only $A_{1/2}$ and $S_{1/2}$ contributions, while resonances with $J\geq 3/2$ can also be excited through $A_{3/2}$ contributions. There are predictions for the $Q^{2}$ dependence of the $A_{1/2}$ and $A_{3/2}$ amplitudes in asymptotic QCD:
@@
where $E_{1+}$, $S_{1+}$, and $M_{1+}$ are the electromagnetic transition multipoles at the mass of the $\Delta(1232)\frac{3}{2}^{+}$ resonance.
@@ 2.4.2   Resonance analysis tools
Figure 2.5: JLab/Hall A data for $\vec{e}p\to e\vec{p}\pi^{0}$ response functions at W=1.232 GeV and $Q^{2}=1.0$ [489]. Notations refer to transverse (t), normal (n) and longitudinal (l) components of the proton recoil polarization. The curves correspond to results obtained using SAID (short dashed), MAID (dashed-dotted), and the dynamical models DMT [377] (dotted), and SL [430] (long-dashed/green). The other curves correspond to Legendre and multipole fits performed by the authors.
Exclusive measurements would in general require full $4\pi$ coverage for the hadronic final state. The only measurement that could claim to be complete was carried out at Jefferson Lab in Hall A [489] employing a limited kinematics centered at resonance for $\vec{e}p\to e^{\prime}\vec{p}\pi^{0}$ at $W=1.232$ GeV, and $Q^{2}\approx 1$ GeV2. Figure 2.5 shows the 16 response functions extracted from this measurement. The results of this measurement in terms of the magnetic $N\Delta$ transition form factor and the quadrupole ratios are included in Fig. 5.5 in Section 5 among other data. They coincide very well with results of other experiments [320, 332, 342, 223] using different analysis techniques that may be also applied to broader kinematic conditions, especially higher-mass resonances. Details of the latter are discussed in  [15, 490]. A similar analysis has recently been carried out within the Jülich-Bonn-Washington model for pion electroproduction multipoles [446], and for combined pion and eta electroproduction [447]. Their results for pion production at the $\Delta(1232)$ compare generally well with the ones used in the analysis [489] but fitted to parameterized amplitudes, while the original analysis was performed at each center-of-mass polar angle separately.
@@
    The Unitary Isobar Model (UIM) was developed in  [376] from the effective Lagrangian approach for pion photoproduction [491]. Background contributions from t-channel $\rho$ and $\omega$ exchanges are introduced and the overall amplitude is unitarized in a K-matrix approximation.
@@
    Dynamical Models have been developed, such as e.g. SAID from pion photoproduction data [492]. The essential feature of the Sato-Lee model developed in [429] is the consistent description of $\pi N$ scattering and pion electroproduction off nucleons. It was utilized in the study of $\Delta(1232)$ excitations in the $ep\to ep\pi^{0}$ channel [430]. The Dubna-Mainz-Taipei model [493] builds in unitarity by directly including the $\pi N$ final state in the T-matrix of photo- and electroproduction.
For some recent reviews of resonance electroproduction and analysis we refer to [18, 15]. The results presented here have been obtained by using both the dispersion relation DR approach as well as the UIM approach. The difference of the two results is the most important contribution to the systematic uncertainty in an evaluation of the model dependence.
@@ 3.1   Historical remarks
It took a long path before all the resonances listed in Tables 2.3 and 2.2 were identified. The RPP 2023 lists 116 light-quark baryons. In the first RPP issue in 1957 [495], no baryon resonance was included, even though the $\Delta^{++}(1232)$ isobar had been observed more than five years earlier. Enrico Fermi had used a $\pi^{+}$ meson beam scattering off protons in a hydrogen target [496]. The cross section showed a sharp rise from the $\pi$ production threshold up to about 1200 MeV. The pion beam’s energy was insufficient to reveal the full resonance profile, but there was a strong indication of the first baryon resonance. In 1961, the PDG recognized the existence of short-lived resonances, reporting the properties of the isospin $I=\nicefrac{{3}}{{2}}$ resonance $\Delta(1232)$ and early evidence for the isospin $I=1$ resonance $\Sigma(1385)$ [497]. The discovery of a double-strange $\Xi(1530)$ [498] as an isospin $I=\nicefrac{{1}}{{2}}$ resonance helped predict an isospin-0 resonance $\Omega^{-}$ at the correct mass [499], paving the way for quarks to be proposed as the fundamental constituents of mesons and baryons [5, 6].
@@
In the quark model, baryons consist of three constituent quarks, each with a spin $s=\nicefrac{{1}}{{2}}$. In $N^{*}$ and $\Delta^{*}$ resonances, all three quarks possess the isospin $I=\nicefrac{{1}}{{2}}$, with the up ($u$) quark having an isospin component $I_{3}=+\nicefrac{{1}}{{2}}$ and charge $Q=+\nicefrac{{2}}{{3}}$ (in units of the elementary charge $e$), while the down ($d$) quark has $I_{3}=-\nicefrac{{1}}{{2}}$ and $Q=-\nicefrac{{1}}{{3}}$. The total quark spin can be $S=\nicefrac{{1}}{{2}}$ or $S=\nicefrac{{3}}{{2}}$, and the total isospin $I=\nicefrac{{1}}{{2}}$ or $I=\nicefrac{{3}}{{2}}$. Baryons with isospin $I=\nicefrac{{1}}{{2}}$ are referred to as nucleons, while $\Delta$ baryons have $I=\nicefrac{{3}}{{2}}$. The $\Delta^{++}(1232)$ corresponds to $|u{\uparrow}u{\uparrow}u{\uparrow}\rangle$, with the three quark spins aligned. This alignment should be forbidden in a ground state with all quarks in an $S$ wave, as the overall wave function would be symmetric, contradicting the Pauli principle. Gell-Mann warned: “Such particles [quarks] presumably are not real, but we may use them in our field theory anyway …" [500]. It took the introduction of para-Fermi statistics [501], later coined “color" by Gell-Mann, to make the overall wave function anti-symmetric. Thus, the $\Delta^{++}(1232)$ resonance may be seen as a harbinger of the development of QCD.
@@
In strange baryons, an up or down quark is replaced by a strange ($s$) quark with isospin $I=0$ and charge $Q=-\nicefrac{{1}}{{3}}$. In cascade baryons, two strange quarks are combined with an up or down quark. The $\Xi^{0}$ and $\Xi^{-}$ baryons decay mostly in a cascade into $\Lambda\pi\to N\pi\pi$. The strangeness quantum number $S=-1$ is assigned to $\Sigma$, $S=-2$ to $\Xi$ baryons. Later, we will shortly discuss properties of baryons with charm and with bottomness, with a charm or bottom quark, $c$ with $Q=\nicefrac{{2}}{{3}}$ and $b$ with $Q=-\nicefrac{{1}}{{3}}$. Top ($t$) quarks are presumably too short-lived to form baryons or mesons, although recent experimental results have sparked a new debate on this matter [502].
@@
The isospin, and thus $SU(2)$ symmetry, was introduced 1932 by Heisenberg [503]. Ne’eman and Gell-Mann [504, 505] organized mesons and baryons in a new symmetry group, $SU(3)$, today called flavor $SU(3)_{f}$, see the inset in Fig. 2.2. Based on this scheme, Gell-Mann and Okubo [505, 506] derived a mass formula, which for baryons can be cast into the form
@@
The relation $\mu_{p}=e/(2m_{q})$ immediately explains why the $g$ factor of the proton is not 2 (or slightly larger) but 5.586 (known with many more digits!). The relevant mass is not the proton mass but the quark mass. Scaling the masses with the $g$ factor, one can estimate the quark masses to $m_{u,d}\approx 340\,{\rm MeV}$. The ratio $\mu_{n}/\mu_{p}$ can be compared to the experimental value $-0.685$. Given this agreement, it came as a surprise when the EMC Collaboration reported that quarks contribute only little to the proton’s spin [507]. For many years, this question remained a topical issue. The spin of quarks and gluons and intrinsic orbital angular momenta all contribute significantly to the proton’s spin. New experiments are planned at several laboratories to improve the accuracy of these determinations, see e.g. [508] for a review.
Back to history: At the Les Houches Summer School in 1965, Dalitz reported on Quark models for the ‘elementary particles’ and sketched the pattern of negative- and positive-parity excitations [509]. More details were presented one year later at the Oxford International Conference on Elementary Particles [510]. For the first time, an interpretation was given within the quark model for the (so far few) nucleon and hyperon resonances. In 1972, color was identified as the source of the strong interaction [511]. The $SU(3)$ structure of color, together with the requirement that the interaction be invariant under local gauge transformations, defined Quantum Chromodynamics (QCD) as the theory of strong interactions [512], with “asymptotic freedom" and “infrared slavery" [513, 514]. The discovery of $J/\psi$ at Brookhaven [515] and SLAC [516] in 1974, the observation of the $\psi^{\prime}$ at SLAC [517], and the interpretation of these narrow resonances using the Cornell potential [518],

### v4

@@ 2.3  Present status
The efforts at the various laboratories discussed above have significantly increased the number of known excitations of the nucleon and their decay modes. A survey of the known light-quark baryons including hyperons is shown in Fig. 2.2. Table 2.3 compares the current status with the status as of 20 years ago ParticleDataGroup:2004fcd . In the mass region from 1850 MeV to 2150 MeV, we now recognize 11 $N^{*}$ resonances, compared to four known in 2004. These 11 resonances have a total overall star rating of 30 stars; in 2004, they had 7 overall stars. The star rating was attributed to their $N\pi$ coupling, with only 11 additional stars given for non-$N\pi$ decay modes, all with one-star evidence. Currently, for resonances in the mass range from 1850 MeV to 2150 MeV, 107 stars are attributed to non-$N\pi$ decay modes. Our knowledge of high-mass resonances comes primarily from elastic scattering of $\pi N$. Table 2.2 lists their PDG masses, widths and uncertainties given in the RPP. For 1* and 2* resonances, the uncertainties are from the measurements; for multiple measurements, the uncertainties span the range of reported data.
Table 2.3: $N^{\ast}$ and $\Delta^{*}$ resonances. Pole positions in terms of the mass $M$ and width $\Gamma$ (in MeV) are taken from the BnGa analysis CLAS:2024iir , Sarantsev:2025lik . Their decay channels as of 2004 ParticleDataGroup:2004fcd are shown in black and those from 2024 in red and in brown. Recent findings are highlighted in blue. Blue stars are suggestions by us. A candidate reported as observed in electroproduction is shown in green Mokeev:2020hhu . Resonances in cyan are not seen in Ref. CLAS:2024iir , Sarantsev:2025lik . Their pole positions are from Ref. ParticleDataGroup:2024cfk . We use the symbol $\sigma$ for the $f_{0}(500)$ throughout this paper.
@@
Figure 2.3: (Color online) Modeling the transition of the universe from the quark-gluon phase to the hadron phase during the first microseconds after the Big Bang. Left: Ratio of strangeness to baryon chemical potential near the critical point in the transition. The shaded area shows the projections from Lattice QCD. The dashed lines represent the hadron resonance gas model projections with all known hadrons as cataloged by the PDG in the RPP editions in 2012 (red dashed-dotted) and 2016 (green dashed-dotted), including ∗∗∗ and ∗∗∗∗ states. Right: The $N^{*}$ and $\Delta^{*}$ resonances are the additional 2016 states with only and ∗∗ states, which are unconfirmed. If a number of those could be confirmed experimentally they could have a significant impact on a more complete understanding of this fundamental hadron phase transition, as the black dashed line indicates. The graphics is adapted from  Chatterjee:2017yhp .
@@ 2.4  Electroproduction of pseudoscalar mesons
The analysis tools for pion and photoproduction of mesons off nucleons are described in detail in recent review articles, in particular in Thiel:2022xtb . The formalism used to analyze the electroproduction data is accessible only in the original articles. Therefore, the methods used are described in some detail in this review.
@@ 2.4.1  Formalism
The kinematics for single $\pi^{+}$ production is shown in Fig. 2.4. The observables of the process $\gamma^{\ast}p\to\pi N^{\prime}$ are often expressed in terms of six helicity amplitudes Aznauryan:2011qj , Walker:1968xu , Berends:1967vi
${H_{i}=\left<\lambda_{\pi};\lambda_{N}|T|\lambda_{\gamma^{\star}};\lambda_{p}\right>,}\vskip-5.69054pt$ (2.6)
where $\lambda$ denotes the helicity of the respective particle, $\lambda_{\pi}=0$, $\lambda_{N}=\pm\frac{1}{2}$, $\lambda_{\gamma^{\star}}=\pm 1,0$, and $\lambda_{p}=\pm\frac{1}{2}$, and $H_{i}$ are complex functions of $Q^{2}$, $W$, and $\theta^{*}_{\pi}$.
@@
where $|{\mbox{$p$}}_{\pi}|$ is the pion three-momentum in the hadronic center-of-mass system, and $K$ is the equivalent real photon lab energy needed to generate a state with mass $W$:
@@
[…]
@@
where $\mbox{$Q$}$$^{*}$ is the photon 3-momentum in the hadronic rest frame. Electromagnetic multipoles are often used to describe the transition from the nucleon ground state to the $\Delta(1232)$, which is predominantly a magnetic dipole transition $M_{1+}$. The electromagnetic multipoles as well as the partial-wave helicity elements are complex quantities and contain both non-resonant and resonant contributions. In order to compare the results to model predictions and lattice QCD, an additional analysis must be performed to separate the resonant parts $\hat{A}_{\pm}$, $\hat{B}_{\pm}$, etc., from the non-resonant parts of the amplitudes. In a final step, the known hadronic properties of a given resonance can be used to determine photocoupling helicity amplitudes that characterize the electromagnetic vertex,
[…]
where the $C^{I}_{\pi N}$ are isospin coefficients. The subscripts $1/2$ and $3/2$ represent the total helicity $\pm 1/2$ and $\pm 3/2$ of the photon-proton system $\gamma^{*}p$ in the initial state. The total transverse absorption cross section for the transition into a specific resonance is given by
@@
and similarly for the longitudinal (scalar) total cross section $\sigma_{L}$. The quantities $A_{1/2},~A_{3/2},~S_{1/2}$ are the transverse and scalar transition amplitudes for resonances excited through photon-proton interactions. Resonances with J=1/2J={1/2} have only $A_{1/2}$ and $S_{1/2}$ contributions, while resonances with $J\geq 3/2$ can also be excited through $A_{3/2}$ contributions. There are predictions for the $Q^{2}$ dependence of the $A_{1/2}$ and $A_{3/2}$ amplitudes in asymptotic QCD:
@@
where $E_{1+}$, $S_{1+}$, and $M_{1+}$ are the electromagnetic transition multipoles at the mass of the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ resonance.
@@ 2.4.2  Resonance analysis tools
Figure 2.5: JLab/Hall A data for $\vec{e}p\to e\vec{p}\pi^{0}$ response functions at W=1.232 GeV and $Q^{2}=1.0$ Kelly:2005jy . Notations refer to transverse (t), normal (n) and longitudinal (l) components of the proton recoil polarization. The curves correspond to results obtained using SAID (short dashed), MAID (dashed-dotted), and the dynamical models DMT Kamalov:2000en (dotted), and SL Sato:2000jf (long-dashed/green). The other curves correspond to Legendre and multipole fits performed by the authors.
Exclusive measurements would in general require full $4\pi$ coverage for the hadronic final state. The only measurement that could claim to be complete was carried out at Jefferson Lab in Hall A Kelly:2005jy employing a limited kinematics centered at resonance for $\vec{e}p\to e^{\prime}\vec{p}\pi^{0}$ at $W=1.232$ GeV, and $Q^{2}\approx 1$ GeV2. Figure 2.5 shows the 16 response functions extracted from this measurement. The results of this measurement in terms of the magnetic $N\Delta$ transition form factor and the quadrupole ratios are included in Fig. 5.5 in Section 5 among other data. They coincide very well with results of other experiments CLAS:2009ces , CLAS:2006ezq , CLAS:2001cbm , Frolov:1998pw using different analysis techniques that may be also applied to broader kinematic conditions, especially higher-mass resonances. Details of the latter are discussed in  Aznauryan:2011qj , Tiator:2011pw . A similar analysis has recently been carried out within the Jülich-Bonn-Washington model for pion electroproduction multipoles Mai:2021vsw , and for combined pion and eta electroproduction Mai:2021aui . Their results for pion production at the $\Delta(1232)$ compare generally well with the ones used in the analysis Kelly:2005jy but fitted to parameterized amplitudes, while the original analysis was performed at each center-of-mass polar angle separately.
@@
    The Unitary Isobar Model (UIM) was developed in  Drechsel:1998hk from the effective Lagrangian approach for pion photoproduction Peccei:1969sb . Background contributions from $t$-channel $\rho$ and $\omega$ exchanges are introduced and the overall amplitude is unitarized in a K-matrix approximation.
@@
    Dynamical Models have been developed, such as e.g. SAID from pion photoproduction data Arndt:2002xv . The essential feature of the Sato-Lee model developed in Sato:1996gk is the consistent description of $\pi N$ scattering and pion electroproduction off nucleons. It was utilized in the study of $\Delta(1232)$ excitations in the $ep\to ep\pi^{0}$ channel Sato:2000jf . The Dubna-Mainz-Taipei model Kamalov:1999hs builds in unitarity by directly including the $\pi N$ final state in the T-matrix of photo- and electroproduction.
For some recent reviews of resonance electroproduction and analysis we refer to Ramalho:2023hqd , Aznauryan:2011qj . The results presented here have been obtained by using both the dispersion relation DR approach as well as the UIM approach. The difference of the two results is the most important contribution to the systematic uncertainty in an evaluation of the model dependence.
@@ 3.1  Historical remarks
It took a long path before all the resonances listed in Tables 2.3 and 2.2 were identified. The RPP 2023 lists 116 light-quark baryons. In the first RPP issue in 1957 Gell-Mann:1957uuj , no baryon resonance was included, even though the $\Delta^{++}(1232)$ isobar had been observed more than five years earlier. Enrico Fermi had used a $\pi^{+}$ meson beam scattering off protons in a hydrogen target Anderson:1952nw . The cross section showed a sharp rise from the $\pi$ production threshold up to about 1200 MeV. The pion beam’s energy was insufficient to reveal the full resonance profile, but there was a strong indication of the first baryon resonance. In 1961, the PDG recognized the existence of short-lived resonances, reporting the properties of the isospin $I=\nicefrac{{3}}{{2}}$ resonance $\Delta(1232)$ and early evidence for the isospin $I=1$ resonance $\Sigma(1385)$ barkas1961data . The discovery of a double-strange $\Xi(1530)$ Pjerrou:1962oxa as an isospin $I=\nicefrac{{1}}{{2}}$ resonance helped predict an isospin-0 resonance $\Omega^{-}$ at the correct mass Gell-Mann:1962yej , paving the way for quarks to be proposed as the fundamental constituents of mesons and baryons Gell-Mann:1964ewy , Zweig:1964jf .
@@
In the quark model, baryons consist of three constituent quarks, each with a spin $s=\nicefrac{{1}}{{2}}$. In $N^{*}$ and $\Delta^{*}$ resonances, all three quarks possess the isospin $I=\nicefrac{{1}}{{2}}$, with the up ($u$) quark having an isospin component $I_{3}=+\nicefrac{{1}}{{2}}$ and charge $Q=+\nicefrac{{2}}{{3}}$ (in units of the elementary charge $e$), while the down ($d$) quark has $I_{3}=-\nicefrac{{1}}{{2}}$ and $Q=-\nicefrac{{1}}{{3}}$. The total quark spin can be $S=\nicefrac{{1}}{{2}}$ or $S=\nicefrac{{3}}{{2}}$, and the total isospin $I=\nicefrac{{1}}{{2}}$ or $I=\nicefrac{{3}}{{2}}$. Baryons with isospin $I=\nicefrac{{1}}{{2}}$ are referred to as nucleons, while $\Delta$ baryons have $I=\nicefrac{{3}}{{2}}$. The $\Delta^{++}(1232)$ corresponds to $|u{\uparrow}u{\uparrow}u{\uparrow}\rangle$, with the three quark spins aligned. This alignment should be forbidden in a ground state with all quarks in an $S$ wave, as the overall wave function would be symmetric, contradicting the Pauli principle. Gell-Mann warned: “Such particles [quarks] presumably are not real, but we may use them in our field theory anyway …" Gell-Mann:1964hhf . It took the introduction of para-Fermi statistics Greenberg:1964pe , later coined “color" by Gell-Mann, to make the overall wave function anti-symmetric. Thus, the $\Delta^{++}(1232)$ resonance may be seen as a harbinger of the development of QCD.
@@
In strange baryons, an up or down quark is replaced by a strange ($s$) quark with isospin $I=0$ and charge $Q=-\nicefrac{{1}}{{3}}$. In cascade baryons, two strange quarks are combined with an up or down quark. The $\Xi^{0}$ and $\Xi^{-}$ baryons decay mostly in a cascade into $\Lambda\pi\to N\pi\pi$. The strangeness quantum number S=1S=-1 is assigned to $\Sigma$, S=2S=-2 to $\Xi$ baryons. Later, we will shortly discuss properties of baryons with charm and with bottomness, with a charm or bottom quark, $c$ with $Q=\nicefrac{{2}}{{3}}$ and $b$ with $Q=-\nicefrac{{1}}{{3}}$. Top ($t$) quarks are presumably too short-lived to form baryons or mesons, although recent experimental results have sparked a new debate on this matter CMS:2025kzt .
@@
The isospin, and thus $SU(2)$ symmetry, was introduced 1932 by Heisenberg Heisenberg:1932aaa . Ne’eman and Gell-Mann Neeman:1961jhl , Gell-Mann:1961omu organized mesons and baryons in a new symmetry group, $SU(3)$, today called flavor $SU(3)_{f}$, see the inset in Fig. 2.2. Based on this scheme, Gell-Mann and Okubo Gell-Mann:1961omu , Okubo:1961jc derived a mass formula, which for baryons can be cast into the form
@@
The relation $\mu_{p}=e/(2m_{q})$ immediately explains why the $g$ factor of the proton is not 2 (or slightly larger) but 5.586 (known with many more digits!). The relevant mass is not the proton mass but the quark mass. Scaling the masses with the $g$ factor, one can estimate the quark masses to $m_{u,d}\approx 340\,{\rm MeV}$. The ratio $\mu_{n}/\mu_{p}$ can be compared to the experimental value 0.685-0.685. Given this agreement, it came as a surprise when the EMC Collaboration reported that quarks contribute only little to the proton’s spin EuropeanMuon:1987isl . For many years, this question remained a topical issue. The spin of quarks and gluons and intrinsic orbital angular momenta all contribute significantly to the proton’s spin. New experiments are planned at several laboratories to improve the accuracy of these determinations, see e.g. Ji:2020ena for a review.
Back to history: At the Les Houches Summer School in 1965, Dalitz reported on Quark models for the ‘elementary particles’ and sketched the pattern of negative- and positive-parity excitations Dalitz:1965fb . More details were presented one year later at the Oxford International Conference on Elementary Particles Dalitz:1966fd . For the first time, an interpretation was given within the quark model for the (so far few) nucleon and hyperon resonances. In 1972, color charge was identified as the source of the strong interaction Fritzsch:1972jv . The $SU(3)$ structure of color, together with the requirement that the interaction be invariant under local gauge transformations, defined Quantum Chromodynamics (QCD) as the theory of strong interactions Fritzsch:1973pi , with “asymptotic freedom" and “infrared slavery" Gross:1973id , Politzer:1973fx . The discovery of $J/\psi$ at Brookhaven E598:1974sol and SLAC SLAC-SP-017:1974ind in 1974, the observation of the $\psi^{\prime}$ at SLAC Abrams:1974yy , and the interpretation of these narrow resonances using the Cornell potential Eichten:1974af ,